OpenStudy (anonymous):

What is the solution set for |-4x - 4| - 28 = -4?

OpenStudy (lgbasallote):

put 28 on the other side first |-4x - 4| = 28 - 4 |-4x - 4| = 24 now solve for x when: -4x - 4 = 24 AND -4x - 4 = -24 got it??

OpenStudy (anonymous):

Yeah I understand it now thank you !

OpenStudy (lgbasallote):

great :D